40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 321 of 325  Not logged in ELOG logo
ID Date Author Type Category Subject
  212   Sat Dec 22 15:32:11 2007 tobinAoGEnvironmentants
Ants are everywhere: on the PSL table, on the circuit board I'm soldering...

I believe I have discovered their energy source.
Attachment 1: DSC_0361.JPG
DSC_0361.JPG
  211   Sat Dec 22 00:52:57 2007 tobinConfigurationPSLISS surgery
In an attempt to quell oscillations in the (unused) outer loop portion of the ISS, I shorted the "PD+" and "PD-" signals from the (nonexistent) outer-loop diodes, and soldered in 47pf compensation capacitors in C92 and C220. This seems to have eliminated oscillations seen at TP41 and TP42. There's still something amiss at TP30 and maybe TP20. Otherwise, the ISS seems happy. I can turn the gain slider to +15dB without saturation (with the HEPA off), though there seems to be less light on the diode (~3.9V) than a week or two ago.
  210   Fri Dec 21 20:32:25 2007 tobinUpdatePhotosGigE camera
I couldn't resist any longer: I plugged in the Prosilica GC 750 GigE camera and took it for a spin. This is the little CMOS camera which sends out video over gigabit ethernet.

There were no difficulties at all in getting it running. I just plugged in the power, plugged in ethernet, and put on a lens from Steve's collection. I downloaded the "Sample Viewer" from the Prosilica website and it worked immediately.

It turns out that "Kirk's" computer has not only a gigabit ethernet card, but a little gigabit ethernet switch. I plugged the camera into this switch. The frame rate is amazing. With the camera under fluorescent lights I thought I saw some wacky automatic gain control, but I think this ~10Hz flicker is aliasing of the 60 Hz room lighting.

I put the camera on the PSL table briefly and tried viewing the image from a laptop over the (54mbs) wireless network. This didn't work so well: you could get a couple frames out of the camera, but then the client software would complain that it had lost communications. It appeared that scattered 1064nm light did show up brightly on the camera image. There is a green ethernet cable currently stashed on the roof of the PSL that appears unused. We can try mounting the gigE CMOS cable in place of one of the CCD video cameras.

I did not try the Linux software.

The camera is currently set up at Kirk's desk, using the cool little tripod Rana got from CyberGuys.

This camera looks very promising! Also, in the test image attached below, a very unusual condition has been documented.
Attachment 1: robs_desk.png
robs_desk.png
  209   Thu Dec 20 21:58:28 2007 AndreySummaryComputer Scripts / ProgramsResults for 2 previous XARM measurements have been added

I attached results (plots) of yesterday's daytime and overnight measurements to the initial reports about those measurements.

These are ELOG entries # 201 and # 205.
  208   Thu Dec 20 21:57:34 2007 AndreyUpdateComputer Scripts / ProgramsMeasurements in XARM today

Today at 2PM I started a program, it should change the suspension gains in the interval from 1.0 to 3.8 with the step 0.2. Estimated running time is till 3.30AM coming night.

Results will be reported on Friday.

BELOW: ADDITION MADE ON FRIDAY EVENING.

Due to some unforeseen circumstances, I was unable to add results on Friday. I have so far accelerometer spectra only, which I add to this ELOG entry.

I have files with the measurement results, and I will process them after Christmas and add to this ELOG entry. I might not be in the lab on Dec. 24 and 25.
Attachment 1: Accelerom_ETMX.png
Accelerom_ETMX.png
Attachment 2: Accelerom_ITMX.png
Accelerom_ITMX.png
  207   Thu Dec 20 19:10:03 2007 waldmanUpdateOMCStressful reattachment of heater
Photos may follow eventually, but for now here's the rundown. I scraped the heater clean of the thermal epoxy using a clean razor blade. Then I stuffed a small piece of lint free cloth in the OTAS bore and wrapped the OMC in tin foil. With a vacuum sucking directly from the face of the OTAS, I gently scraped the glue off the OTAS aluminum. I wiped both the OTAS and the heater down with an isoproponal soaked lint-free cloth. I put a thin sheen of VacSeal on the face of the heater, wiping off the excess from the edges with a cloth. Then I clamped the heater to the OTAS using 2" c-clamps from the tombstone back to the heater front, making sure the alignment of the OTAS was correct (connector on the absolute bottom, concentric with the OTAS outer diameter). I added a second clamp, then beaded the outside of the joint with a little bit extra VacSeal, just for kicks. I'll leave it covered at least overnight, and maybe for a day or two.

sam
  206   Thu Dec 20 19:05:34 2007 waldmanHowToOMCHOWTO build front ends
For instance, to build the TPT front end code.

  • Save your file /cvs/cds/advLigo/src/epics/simLink/tpt.mdl
  • go to /cvs/cds/advLIGO on the TPT machine
  • do make clean-tpt tpt install-tpt
  • do rm /cvs/cds/caltech/chans/daq/C2TPT.ini (this step is needed because the DAQ install code isn't quite right at the time of this writing.
  • do make install-daq-tpt
  • run starttpt to restart the tpt computer.

Enjoy.
  205   Thu Dec 20 02:04:09 2007 AndreyUpdateComputer Scripts / ProgramsNew overnight measurements in XARM and their results

I ran in the daytime/evening time my program, changing the damping gains in suspension "position" degree of freedom for ETMX and ITMX
in the interval from 1.00 to 3.75 with the step 0.25 (see entry # 201).

Now I am running overnight (from 2AM till 9AM) the program changing the gains in the interval from 1.3 to 3.5 with the step 0.20,
12 X 12 = 144 experimental points. I started so late because I fell asleep after my Wednesday evening dinner, then woke up half an hour ago and hurried to the lab.

BELOW: RESULTS OF MEASUREMENTS WERE ADDED ON THURSDAY EVENING, DEC. 20.

All the meaning of the attachments 1-3, 4-6, 7-9, 10-11 is the same as in previous ELOG entries # 195, # 199, # 202, see in those entries which graph corresponds to which coordinate axes orientation.
Attachment 1: RMS-08Hz-Top-View.png
RMS-08Hz-Top-View.png
Attachment 2: RMS-3Hz-Top-View.png
RMS-3Hz-Top-View.png
Attachment 3: RMS-broadband-Top-View.png
RMS-broadband-Top-View.png
Attachment 4: RMS-08Hz-Side_View.png
RMS-08Hz-Side_View.png
Attachment 5: RMS-3Hz-Side_View.png
RMS-3Hz-Side_View.png
Attachment 6: RMS-broadband-Side_View.png
RMS-broadband-Side_View.png
Attachment 7: RMS-08Hz-Q_I-Q_E-Axes.png
RMS-08Hz-Q_I-Q_E-Axes.png
Attachment 8: RMS-3Hz-Q_I-Q_E-Axes.png
RMS-3Hz-Q_I-Q_E-Axes.png
Attachment 9: RMS-broadband-Q_I-Q_E-Axes.png
RMS-broadband-Q_I-Q_E-Axes.png
Attachment 10: Accelerometer-ETMX.png
Accelerometer-ETMX.png
Attachment 11: Accelerometer-ITMX.png
Accelerometer-ITMX.png
  204   Wed Dec 19 20:28:27 2007 AndreyDAQPEMNames for all 6 accelerometers have been changed

I eventually changed the names for all 6 accelerometers (see my ELOG entry # 172 from Dec. 05 about my intent to do that).

I removed the word "BS" from their names,
and I changed the word combination "ACC_BS_EAST" in the old name for "ACC_ITMX" in the new name;
as well "ACC_BS_WEST" is now replaced by "ACC_ETMX".
(the reasoning behind such a change should become clear from my ELOG entry #172).

New accelerometer names are:
(note: there are no spaces (nowhere!) in the names of accelerometers, but ELOG replaces ": P" written without a space by a strange symbol Tongue)

C1 : PEM - ACC _ ETMX _ X ;
C1 : PEM - ACC _ ETMX _ Y ;
C1 : PEM - ACC _ ETMX _ Z ;
C1 : PEM - ACC _ ITMX _ X ;
C1 : PEM - ACC _ ITMX _ Y ;
C1 : PEM - ACC _ ITMX _ Z .

One can find them in "C1 : PEM - ACC" in Dataviewer.

  203   Wed Dec 19 16:40:12 2007 steveUpdateSAFETYlaser safety glasses measured
I measured the coarse transission at 1064nm of the 40m safety glasses today.

12 pieces of UVEX # LOTG-YAG/CO2 light green, all plastic construction, ADSORBANT

3 pieces of 6KG5, Scott colored filter glass type,

individual prescription glasses: alan, bob, ben, jay and steve

7 pieces of dual waveleght glasses

These glasses showed 0.00mW transmission out of 170mW Crysta Laser 1064
  202   Wed Dec 19 16:07:37 2007 AndreySummaryComputer Scripts / ProgramsResults of overnight measurements Tue/Wed night (entry #198)

As indicated in ELOG entry 198, I was making overnight measurements during last night from Tuesday to Wednesday.

I was changing the suspension damping gain in ETMX and ITMX in "position" degree of freedom between values of 1.00 and 4.50 with the step 0.25.

Results for RMS of peaks (A) at 0.8Hz, (B) at about 3.0Hz and (C) in the range from 0.6Hz to 3.7Hz ("RMS in a broad interval") are given below:

I plotted three results for RMS in the abovementioned three intervals in three different ways:

1) view from the top in the axes (Q_{ITMX}+Q_{ETMX})/2 and (Q_{ITMX}-Q_{ETMX}) -> first three graphs (attachments 1 -3);

2) view from the side in the same sum- and difference-axes -> next three graphs (attachments 4-6);

3) view from the side in Q_{ITMX} and Q_{ETMX} axes -> next three graphs (attachments 7-9)

Attachments 10 and 11 show ratios of accelerometer signals at different times of the night/morning.


A little discussion about these graphs:

1) The areas of minima and of rapid growth are the same for all the measurements during all three nights.

2) Tonight there was a strange spike for the values of Q_{ETMX}=2.5 and Q_{ITMX}=4.0. I interpret that as an error of experiment.

3) On all the plots from all three nights there is a wide area of minimum on the plots for RMS at 0.8Hz and for "RMS in the broad interval",
and the graph for "RMS at 3Hz" indicates a clearer minimum in a localized area for Q_{ITMX}=2+-1, Q_{ETMX}=2+-1. Note that this area 2+-1
is included into the wide region of minimum for "RMS at 0.8Hz" and "RMS in a broad range".

Therefore, my guess at this stage is that we can choose the optimized value of suspension damping gains for both Q_{ITMX} and Q_{ETMX} somewhere
around 2+-1. I would like to make another overnight measurement (tonight) in that narrowed region with a small step to have more certainty.

By the way, I realized that I was a little bit careless and at some plots Q_I stands for {Q_ITMX}, and Q_E stands for Q_{ETMX}.
Attachment 1: RMS_08Hz_Top_view.png
RMS_08Hz_Top_view.png
Attachment 2: RMS_3Hz_Top_view.png
RMS_3Hz_Top_view.png
Attachment 3: RMS_broad_Top_view.png
RMS_broad_Top_view.png
Attachment 4: RMS_08Hz_Side_view.png
RMS_08Hz_Side_view.png
Attachment 5: RMS_3Hz_Side_view.png
RMS_3Hz_Side_view.png
Attachment 6: RMS_broadband_Side_view.png
RMS_broadband_Side_view.png
Attachment 7: RMS_08Hz_Q_I-Q_E-axes.png
RMS_08Hz_Q_I-Q_E-axes.png
Attachment 8: RMS_3Hz_Q_I-Q_E-axes.png
RMS_3Hz_Q_I-Q_E-axes.png
Attachment 9: RMS_broadband_Q_I-Q_E-axes.png
RMS_broadband_Q_I-Q_E-axes.png
Attachment 10: Accelerom_ETMX.png
Accelerom_ETMX.png
Attachment 11: Accelerom_ITMX.png
Accelerom_ITMX.png
  201   Wed Dec 19 15:51:00 2007 AndreyUpdateComputer Scripts / ProgramsDaytime measurements in XARM and their results

I was making measurements in XARM for three different nights. All the results agree with each other (I will put the results from the last night soon).

Steve Vass recommended to me to compare those results with the daytime data, in order to see if there is a real necessity to run the scripts overnight or if daytime results will yield similar results.

XARM has been locked, and I am taking measurements today from 3.30PM till 11.30PM.

I will be changing the suspension damping gains in ETMX and ITMX "position" degrees of freedom in the interval from 1.0 to 3.75 with the step 0.25.

BELOW: RESULTS OF MEASUREMENTS WERE ADDED ON THURSDAY, DEC. 20.

All the meaning of the attachments 1-3, 4-6, 7-9, 10-11 is the same as in previous ELOG entries # 195, # 199, # 202, see in those entries which graph corresponds to which coordinate axes orientation.
Attachment 1: RMS-08Hz-Top_View.png
RMS-08Hz-Top_View.png
Attachment 2: RMS-3Hz-Top_View.png
RMS-3Hz-Top_View.png
Attachment 3: RMS-broadband-Top_View.png
RMS-broadband-Top_View.png
Attachment 4: RMS-08Hz-Side-View.png
RMS-08Hz-Side-View.png
Attachment 5: RMS-3Hz-Side_View.png
RMS-3Hz-Side_View.png
Attachment 6: RMS-broadband-Side_View.png
RMS-broadband-Side_View.png
Attachment 7: RMS-08Hz-Q_I-Q_E-Axes.png
RMS-08Hz-Q_I-Q_E-Axes.png
Attachment 8: RMS-3Hz-Side_View.png
RMS-3Hz-Side_View.png
Attachment 9: RMS-broadband-Side_View.png
RMS-broadband-Side_View.png
Attachment 10: Accelerometer_ETMX.png
Accelerometer_ETMX.png
Attachment 11: Accelerometer_ITMX.png
Accelerometer_ITMX.png
  200   Wed Dec 19 11:31:01 2007 steveOmnistructurePEMaircond filter maintenance
Jeff is working on all air condiontion units of the 40m lab
This we do every six months.
Attachment 1: acfilters6m.jpg
acfilters6m.jpg
  199   Tue Dec 18 23:41:00 2007 AndreySummaryComputer Scripts / ProgramsResults of Mon/Tue overnight measurements (entry #194)

Here I inform our community about the results of the measurements of RMS values in XARM during the previous night from Monday to Tuesday (I announced those measurements in ELOG entry #194).

All the plots in today's report seem to agree well with the analogous plots from the night from Saturday to Sunday (those results are given in ELOG entry # 195).

All the intervals in which RMS have been calculated are the same as in yesterday's ELOG entry #195.

I plotted three results for RMS in the abovementioned three intervals in three different ways:

1) view from the top in the axes (Q_{ITMX}+Q_{ETMX})/2 and (Q_{ITMX}-Q_{ETMX}) -> first three graphs (attachments 1 -3);

2) view from the side in the same sum- and difference-axes -> next three graphs (attachments 4-6);

3) view from the side in Q_{ITMX} and Q_{ETMX} axes -> next three graphs (attachments 7-9, also attch. 12), above accelerometer spectra (attachments 10-11).

Also, I compared the ground noise level by comparing spectra of accelerometer signals at different times during that night. As a reminder, before my disease I installed one accelerometer near ITMX and another accelerometer near ETMX (see entries 161 and 172 in ELOG). The plots of ratios of accelerometer signals at different times (pairs of times that were used: 11PM and 2AM, 11PM and 5AM, 11PM and 8AM) are given below, see attachments 10-11. The program was running from 11PM on Monday till 9AM on Tuesday.

As I explained in the previous ELOG entry # 198, tonight I am taking experimental data in the narrowere interval from 1.00 to 4.50 with a smaller step 0.25.
Attachment 1: RMS_08HZ_Top_View.png
RMS_08HZ_Top_View.png
Attachment 2: RMS_3HZ_Top_View.png
RMS_3HZ_Top_View.png
Attachment 3: RMS_broad_Top_View.png
RMS_broad_Top_View.png
Attachment 4: RMS_08HZ_Side_View.png
RMS_08HZ_Side_View.png
Attachment 5: RMS_3HZ_Side_View.png
RMS_3HZ_Side_View.png
Attachment 6: RMS_broad_Side_View.png
RMS_broad_Side_View.png
Attachment 7: RMS_08HZ_Q_E_Q_I_Axes.png
RMS_08HZ_Q_E_Q_I_Axes.png
Attachment 8: RMS_3HZ_Q_E_Q_I_Axes.png
RMS_3HZ_Q_E_Q_I_Axes.png
Attachment 9: RMS_broad_Q_E_Q_I_Axes.png
RMS_broad_Q_E_Q_I_Axes.png
Attachment 10: Accelerometer_ITMX.png
Accelerometer_ITMX.png
Attachment 11: Accelerometer_ETMX.png
Accelerometer_ETMX.png
Attachment 12: RMS_broad_Q_E_Q_I_Axes.png
RMS_broad_Q_E_Q_I_Axes.png
  198   Tue Dec 18 23:27:36 2007 AndreyConfigurationComputer Scripts / ProgramsNew overnight measurements (this night from Tue to Wed)

I am making overnight measurements in XARM tonight.

This is the third night of measurements in XARM, but tonight I am scanning the narrower region between values of damping gain 1.00 and 4.50 with the smaller step 0.25. (for comparison, during two previous measurements the region was between 1.0 and 7.0 with the step 0.5).

I have relocked the XARM before the start of the measurements.

I started running the program at 9.30PM, and it should collect all the data by 9.00AM wednesday morning.

Below are explanations why I chose these different parameters for the interval and step:

I am going to put the results of previous night measurements into the next ELOG entry, and it we be pretty obvious from those graphs that results in XARM from the two previous (different) nights agree well with each other, and the approximate positions of minima and areas of "big growth" of the surfaces are pretty obvious from those graphs. It is clear that RMS are too big for the values of the damping gain bigger than 4.0, and that minima are somewhere near the values of 2.0. But those graphs were too rough to locate a somewhat precise value for the minima. Therefore, I am studying tonight the interval of gains between 1.00 and 4.50 with a smaller step.

A short note how I estimate time that is necessary to collect the experimental data.

there are 15 experimental points for each ETMX and ITMX suspension gains in the interval between 1.00 and 4.50 with the step 0.25. They are: 1.00, 1.25, 1.50, 1.75, 2.00, ..., 3.75, 4.00, 4.25, 4.50. As I am changing both ETMX and ITMX gains, I have an array of 15*15=225 elements.
It takes 3 minutes for each point to collect the data (I wrote the program that way). Therefore, the total time it takes to run the program is: 225*3=675 minutes, or 675/60=11.25 hours, almost 11 and a half hours.
  197   Tue Dec 18 21:31:31 2007 tobinUpdatePSLISS RIN
My measurements of the ISS RIN via the SR785 and via the DAQ disagree considerably. The spectral shapes are very similar, however, so I expect that a constant factor is creeping in somewhere. Measurements taken at the PD DC monitor points using the SR785 attached. There is a lot of excess noise in the 300 Hz - 1 kHz region.
Attachment 1: iss-rin.pdf
iss-rin.pdf
  196   Tue Dec 18 16:50:35 2007 tobinUpdateSAFETYuvex laser safety glasses defective
A few days ago we noticed what appeared to be a blotched, speckled fracturing of the coating of the "UVEX" laser safety glasses. These are the glasses with "transparent" (reflective to 1064nm) lenses and white frames that we keep in a box on top of a filing cabinet in the control room. Today Steve measured the transmission of these glasses and found 80% transmission of 1064nm in several cases.

Do not use the white, transparent "uvex" laser safety glasses until further notice. Steve has hidden them away so that you won't be tempted.

Below is attached a photo of a bad lens.
Attachment 1: bad-glasses.jpg
bad-glasses.jpg
Attachment 2: bad-glasses-zoom.jpg
bad-glasses-zoom.jpg
  195   Tue Dec 18 00:51:39 2007 AndreyUpdateComputer Scripts / ProgramsResults of Saturday overnight measurements

As I indicated in the previous e-log entry (#194), I made overnight measurements in XARM in the night from Saturday to Sunday.

Root-mean-square values of the peaks in calibrated spectra were calculated, and I plotted them as functions of suspension gains in ITMX and ETMX "position" degrees of freedom.
More specifically, Q_ITMX means the value in the channel "C1:SUS-ITMX_SUSPOS_GAIN", while Q_ETMX means the value in the channel "C1:SUS-ETMX_SUSPOS_GAIN".

Root-mean-square values (RMS) were calculated during that night in three intervals:

1) around 0.8 HZ in the interval (0.6 Hz <-> 1.0 Hz);

2) around 3.0 Hz in the interval (2.0 Hz <-> 3.6 Hz);

3) in the broad interval from 0.6Hz to 3.6Hz.


I plotted three results for RMS in the abovementioned three intervals in three different ways:

1) view from the top in the axes (Q_{ITMX}+Q_{ETMX})/2 and (Q_{ITMX}-Q_{ETMX}) -> first three graphs (attachments 1 -3);

2) view from the side in the same sum- and difference-axes -> next three graphs (attachments 4-6);

3) view from the side in Q_{ITMX} and Q_{ETMX} axes -> next three graphs (attachments 7-9), above accelerometer spectra (attachments 10-11).


Also, I compared the ground noise level by comparing spectra of accelerometer signals at different times during that night. As a reminder, before my disease I installed one accelerometer near ITMX and another accelerometer near ETMX (see entries 161 and 172 in ELOG). The plots of ratios of accelerometer signals at different times (pairs of times that were used: 12AM and 3AM, 12AM and 6AM, 12AM and 9AM) are given below, see attachments 10-11.

Tomorrow I will try to compare the results with the second measurements that are being taken tonight.
Attachment 1: RMS_08Hz_top_view.png
RMS_08Hz_top_view.png
Attachment 2: RMS_3Hz_top_view.png
RMS_3Hz_top_view.png
Attachment 3: RMS_broad_top_view.png
RMS_broad_top_view.png
Attachment 4: RMS_08Hz_Qsum-Qdiff-axes.png
RMS_08Hz_Qsum-Qdiff-axes.png
Attachment 5: RMS_3Hz_Qsum-Qdiff-axes.png
RMS_3Hz_Qsum-Qdiff-axes.png
Attachment 6: RMS_broad_Qsum-Qdiff-axes.png
RMS_broad_Qsum-Qdiff-axes.png
Attachment 7: RMS_08Hz_Qaxes.png
RMS_08Hz_Qaxes.png
Attachment 8: RMS_3Hz_Qaxes.png
RMS_3Hz_Qaxes.png
Attachment 9: RMS_broad_Qaxes.png
RMS_broad_Qaxes.png
Attachment 10: Accel_ITMX.png
Accel_ITMX.png
Attachment 11: Accel_ETMX.png
Accel_ETMX.png
  194   Mon Dec 17 23:42:08 2007 AndreyConfigurationComputer Scripts / ProgramsOvernight measurements in X-arm

I am making overnight measurements this night (from Monday to Tuesday) in XARM.

The X-arm is now locked, and the values for suspension damping gain will be changed in the interval from 1 to 7 with the step 0.5 in both ITMX and ETMX.

This is the second, repeated measurement. The results of the first measurement from Saturday to Sunday night will be reported in the separate ELOG entry (sorry, I did not make an ELOG entry on Saturday evening about running the program overnight).

The very first attempt to run the script in the night from Thursday to Friday was not successful.
  193   Mon Dec 17 11:47:13 2007 albertoUpdateElectronicsan alternative design for the RFAM monitor's filter at 33Mhz
Since the Butterworth turned out o be rather wide-band, I tried an other configuration for the 33 MHz filter. Attached are the simulated transfer function and the measured. As one can see, the measured peak is much broader than expected.
Attachment 1: RFSim99-33MHz.png
RFSim99-33MHz.png
Attachment 2: RF99-SimmButterworthPrototype.png
RF99-SimmButterworthPrototype.png
Attachment 3: RFSim99-33MHz-TFplot.png
RFSim99-33MHz-TFplot.png
  192   Sun Dec 16 16:52:40 2007 dmassUpdateComputersComputer on the end Fixed
I had Mike Pedraza look at the computer on the end (tag c21256). It was running funny, and turns out it was a bad HD.

I backed up the SURF files as attachments to their wiki entries. Nothing else seemed important so the drive was (presumably) swapped, and a clean copy of xp pro was installed. The username/login is the standard one.

Also - that small corner of desk space is now clean, and it would be lovely if it stayed that way.
  191   Thu Dec 13 23:56:02 2007 AndreyConfigurationComputer Scripts / ProgramsOvernight measurements

After my disease (fever, vomitting, nose problem, overall weakness) I returned to LIGO today for the first time after the weekend, and I am running the script for the XARM-measurements over this night.

So, suspension dumping gains should undergo changes in the interval from 1 to 10 in both ITMX and ETMX.

XARM has been of course locked.

I started running the script for the first time at about 10PM, but I realized after an hour and a half that my step of gain increase 0.2 was too shallow, too small to execute my program during one night. Therefore, I needed to terminate the program, change my program so that it increases the gain with increment 0.5, not 0.2, and started it again around midnight.

Going home.

P.S. The red pump that I borrowed from the lab (Steve's pump?) is back at its previous place. The tire-worker tells me that I absolutely need to change all four tires for almost 500 dollars. I regret a lot about that huge money loss.
  190   Thu Dec 13 12:05:36 2007 albertoOmnistructureElectronicsThe new Butterworth seems to work quite well
It works better probably because of the small inductors I'm using this time.
The peak is at 30 MHz because I didn't have the precise elements to get 33.

The bandwidth and the Q could be improved by adding one or two more order to the filter and trying to better match the low-pass' resonant frequency with the high-pass'.

Also I have to see if it could work at 166 and 199 MHz as well.
Attachment 1: TF_New_Butterworth_12-Nov-2007_TF.png
TF_New_Butterworth_12-Nov-2007_TF.png
Attachment 2: Bultervverth2.png
Bultervverth2.png
  189   Wed Dec 12 22:24:48 2007 tobinFrogsPEMweather station
I poked at the weather station briefly this evening.

* There's almost nothing in the elog about it.
* It exists. It is located on the North wall, just north of the beam splitter.
* It seems to be displaying reasonable data for the indoors, but nothing for the outdoor sensors.
* c1pem didn't seem to be starting up (couldn't telnet into it, etc). I altered its startup file and reset it several times, and eventually it came to life.
* the weather station has a serial cable that goes all the way to c1pem. I plugged it in.
* however, the Weather.st program complains "NO COMM"--it gets no data from the weather station
* The next thing to do is to plug in a laptop to that serial cable and see if the weather station can be convinced to talk.
  188   Wed Dec 12 16:22:22 2007 albertoOmnistructureElectronicsLC filter for the RF-AM monitor circuit
In the LC configuration (see attached schematic) the resonant frequency is tuned to one of the peak of our RF-AM monitor and it is amplified by a factor equal to the Q of the filter. As I wrote in one of the last elog entries, we would like amplifications of about 10-30 dB in order to have negligible couplings. Such values are obtained only with small capacitances (few or less pF). The drawback is relatively large inductance (uH or more) which has inevitably low Self Resonant Frequencies (SRF - the resonant frequencies of the RLC circuit usually associated with an actual inductor - ~ MHz). Even before, one limit is also the input impedance of the RF amplifier. Quality factors > 1 require megaohms, far from the 50 ohms in the MiniCircuit amplifiers Iím using now. So, if we plan to use these even for the final design of the circuit, we have to abandon the LC configuration.
For this same reason the only way I could get the expected responses from my several test boards was with a 10 megaohm input probe (see attachment for the measurement with and without probe). Assuming that impedance, I found these as the best trade-offs between the attenuation requirements and the values of the inductors for respectively the peaks at 33, 66,133, 166,199 MHz:
26uH, 6.6u, 20u, 73u, 16u
If we could find inductor with these values and high SRF the configuration should work. The problem is I couldnít find any. Above a few uH they all seem to have SRF ~ MHz.
That is why I switched to the Butterworth. This should work despite the input impedance of the amplifier and with much smaller inductances. I made a totally new test circuit, with surface mount components. I think I still have to fix some things in the measurements but (this time I got rid of the simulator I was using earlier and designed a new configuration with new values from the Horowitzís tables) it seems I have the expected peaks. More soon.
Attachment 1: TF_LC_filter_10pF_1.8uH_scope_probe.png
TF_LC_filter_10pF_1.8uH_scope_probe.png
Attachment 2: TF_LC_filter_10pF_1.8mH_no_probe.png
TF_LC_filter_10pF_1.8mH_no_probe.png
Attachment 3: LC_filter_schematic.png
LC_filter_schematic.png
  187   Mon Dec 10 20:35:59 2007 tobinConfigurationComputer Scripts / Programsautolocking scripts
I added this tidbit of csh code to the MZ autolocker to prevent multiple copies from running (on one computer):
if (`pgrep lockMZ | wc -l` > 1) then
   echo lockMZ is already running! 
   exit
endif
Similarly, here's some bash code that does something similar; I'll add it to the other autolocker scripts:
if                                                                                                                       
  pgrep `basename $0` | grep -v $$ > /dev/null                                                                           
then                                                                                                                     
  echo Another copy of this program is already running.  Exiting!                                                        
  exit 1                                                                                                                 
fi
This code searches for all processes with the same name as this script ($0) and then use grep to exclude (-v) the current process ID ($$).
  186   Mon Dec 10 19:08:03 2007 tobinConfigurationPSLMZ
The MZ seems finicky today--it keeps unlocking and relocking.

I've temporarily blocked one of the MZ arms while I work on the ISS.
  185   Mon Dec 10 18:42:20 2007 tobinUpdatePSLISS RIN script
I wrote a script to measure the ISS RIN. The script uses the "labca" interface (described in an earlier entry) to read and twiddle EPICS settings and mDV to get DAQ data. The script measures open loop RIN, closed loop RIN at each of several gain slider settings, and dark noise. The dark noise is obtained by misaligning (unlocking) the PMC. The script also compares the whitened and unwhitened spectra for an open loop measurement and performs a fit of a simple pole to find the dewhitening filter.

This is all very exciting, but I don't quite believe the results, since the closed loop RIN seems to bottom out at 2e-7/rtHz regardless of the gain slider setting.

Sample output attached. The script may be found at scripts/PSL/ISS/rin.m.
Attachment 1: rin-20071210-1831.pdf
rin-20071210-1831.pdf
  184   Mon Dec 10 13:54:26 2007 robHowToComputer Scripts / ProgramsDon't blame the Matlab compiler

For human error. We should be careful to only run the compiler outside the mDV directory (thus placing the _mcr outside of the range the addpath command in the mdv_config files). Or maybe there's a smarter solution...
  183   Fri Dec 7 19:14:30 2007 tobinUpdatePSLISS dark noise - ground loop enlightenment
My alleged 60 Hz harmonics were all from a ground loop created by connecting the SR785 ground to the ISS circuit ground; they disappeared when I set the SR785 input to "floating ground." doh!

I modified the ISS PD's to have a 100 ohm resistor in series with the output (in place of 20 ohms). The diodes are again in place on the table, ready for action.
  182   Fri Dec 7 18:31:30 2007 tobinUpdateComputer Scripts / Programscompiled matlab hoses itself
Addendnum. The reason the linemon_mcr command was in the path was because of the user issuing the command "addpath(genpath(pwd))" where genpath(D) "returns a path string starting in D, plus, recursively, all the subdirectories of D."

The Matlab compiler is still bad, however.
  181   Fri Dec 7 18:28:30 2007 tobinUpdateComputer Scripts / Programscompiled matlab hoses itself
Andrey pointed out to me that some matlab functions in the Signal Processing Toolbox were dying with errors. Looking into the .m file (identified using the "which" command), I was surprised to see binary garbage rather than glistening, clear Matlab prose. Then I noticed the directory in which it was finding the .m file:
>> which decimate
/cvs/cds/caltech/apps/mDV/extra/linetrack_c_mcr/toolbox/signal/signal/decimate.m
See that "linetrack_c_mcr" directory? This is what is generated when a "compiled" (grumble) Matlab program is run--it decompresses itself into a subdirectory containing weird semi-compiled binary .m files. Unfortunately this is somehow getting incorporated into the matlab path. (I assume there is something in mDV that says "put all subdirectories into the path.")

I hate the Matlab compiler.
  180   Fri Dec 7 14:14:48 2007 robMetaphysicsComputer Scripts / Programstdsread problems on Solaris

tdsread has developed a strange new illness, whereby it cannot read EPICS values from two subsystems at once (e.g., getting an LSC and SUS value simultaneously). I thought this might have something to with the fact that both losepics and iscepics are running on the same box,
but the same thing happens with IOO EPICS records, so that's not the culprit.

This is new behaviour, and it's only happening on the solaris machines. I suspect some ENV/cshrc juju has caused it, as the tdsread executable is the same one from April, and I don't think our EPICS infrastructure has changed otherwise. In the near term we can either try running the scripts on linux, or modify the IFO scripts to not do these types of calls.
  179   Fri Dec 7 11:33:24 2007 waldmanOmnistructureOMCPZT wiring
The 2 pin LEMO connector has got an unmarked pin and a pin marked by a white half-circle.
The unmarked pin is connected to the side of the PZT attached to the mirror.
The marked pin is connected to the side of the PZT attached to the tombstone.
  178   Fri Dec 7 00:02:26 2007 ranaSummaryIOOMC/FSS Frequency Noise
The FSS frequency noise is not very bad.

I compared the MC_F spectra between Hanford and the 40m using DTT and its 'User NDS' option.
After Sam, Jenne, and DavidM installed the new MC Servo some time ago, the MC_F spectrum here
has had some whitening before it goes into the DAQ (on board; same as LLO & LHO). The tuning
coefficient of the VCO is also basically the same between all PSLs since everyone has the same
chip in the VCO driver.

Therefore, at the frequencies where the MC gain is more than ~4, the MC_F signal calibration is
the same here as anywhere. Since its the servo control signal, its basically a measure of the
frequency noise incident on the MC -- its just what comes out of the FSS with the table noise on
top. At low frequencies (< 100 Hz) its a measure of the motion of the MC mirrors.

Above 200 Hz ours is the same as theirs; except for the enormous power line spikes. I think that's
either all on the light. But our acoustics are better and the noise above 1 kHz levels off at the
same flat floor (the phase noise of the VCO) as H1. The huge lump around 100 Hz is the MC2 DAC noise and
it goes down to the H1 levels when we flip on the dewhites. The giant excess from 5-50 Hz is just the fact
that our stacks don't do much until 20-30 Hz.

So we can stop blaming the FSS and move on with life as soon as Tobin gets the ISS back in shape.
Attachment 1: fly.pdf
fly.pdf
  177   Thu Dec 6 19:30:43 2007 tobinUpdatePSLISS dark noise - 60 Hz!
A higher resolution spectrum [attached] shows that nearly all of the excess dark noise on the ISS is in 60 Hz harmonics (with some 256 Hz harmonics too--are these from the DAQ?).

With the loop closed and the slider at 5dB, the laser light coming out has a noise floor of 10^-7 RIN or better from 40 Hz to 8 kHz.

Now to figure out why all this 60 Hz is getting in... (I tried turning off all the lights and the HEPA, and moving the SR785 further away, none of which did anything.)
Attachment 1: iss.pdf
iss.pdf
  176   Thu Dec 6 19:19:47 2007 AndreyConfigurationSUSSuspension damping Gain was restored

Suspension damping gain was disabled for some reason (all the indicators in the most right part of the screen C1SUS_ETMX.adl were red), it is now restored.
  175   Thu Dec 6 18:11:15 2007 robHowToComputer Scripts / ProgramsMaking DMF monitors

I was able to use the matlab compiler to compile a version of the linetracker written by Rana, and run the compiled version on mafalda.

I believe I made the necessary edits to our mDV config file so that it should be easy for others to follow these steps:

1) Write the DMF routine you want, as a matlab function (not a script).

2) If it runs correctly in matlab, then from the matlab command line compile
it using the -m flag (i.e., mcc -m myfun.m). You should run the
compiler from the directory where you want the executable to end up (don't use the mDV/extra
directory so it doesn't get all cluttered).

3) prior to running the resulting executable (which should be called simply myfun),
prepend the directories
/cvs/cds/caltech/apps/linux/matlab/bin/glnx86
/cvs/cds/caltech/apps/linux/matlab/sys/os/glnx86/

to the LD_LIBRARY_PATH enviroment variable. These directories must be prepended as the
versions that already exist in /usr/lib don't work; I'm loathe to do this in the cshrc.40m
for fear of later conflicts, so it will need to be done in some sort of shell script which
calls the matlab executable.
  174   Thu Dec 6 15:22:42 2007 AndreySummaryElectronicsPictures of the inside of He-Ne laser

Steve gave me an old "dead" He-Ne laser that long time ago was used for ETMX optical lever.

I dismantled it (cutting the metallic enclosure with a metallic saw), and these are two pictures of what is inside.
Attachment 1: DSC_0226.JPG
DSC_0226.JPG
Attachment 2: DSC_0228.JPG
DSC_0228.JPG
  173   Thu Dec 6 15:21:59 2007 albertoFrogsElectronicsRF Transfer Function of Stiff Aluminum Wires
Transfer function of 3cm long Aluminum wires and of 3cm stranded wires
Attachment 1: TF_3cm_stiff_wires.amplitude.png
TF_3cm_stiff_wires.amplitude.png
Attachment 2: DSC_0225compressed.JPG
DSC_0225compressed.JPG
Attachment 3: TF_3cm_stranded_wires.amplitude.png
TF_3cm_stranded_wires.amplitude.png
  172   Wed Dec 5 23:19:03 2007 AndreyConfigurationPEMAccelerometers are turned on

All accelerometers have been turned on, as Alan asked during Wednesday meeting.

Typical power spectra and coherence plots are attached below.

"East" in the name means that the previous location of accelerometrs was to the east from "Beamsplitter" (the location for "east" accelerometers was not changed, actually, it is still near ITMX), while "west" means that previously accelerometers were to the west from the BS, but now their new location is near the ETMX.

I will change the names of the channels tomorrow (Thursday) when someone (Tobin?) will show to me how to do it.

P.S. (addition made on Dec. 19th, 2007, by Andrey) I intended to change the names of accelerometers the next day, Thursday Dec. 06,
but I did not do it that day (did not understand how to do it), then I fell ill, and eventually
I changed the names of accelerometers on December 19th, see entry to ELOG #204)
Attachment 1: Power_Sp_and_Coh_XY-EAST.pdf
Power_Sp_and_Coh_XY-EAST.pdf
Attachment 2: Coherence-ZX_East.pdf
Coherence-ZX_East.pdf
Attachment 3: Coherence-ZY_East.pdf
Coherence-ZY_East.pdf
Attachment 4: Power_Sp_WEST.pdf
Power_Sp_WEST.pdf
Attachment 5: Coherence-ZX_West.pdf
Coherence-ZX_West.pdf
Attachment 6: Coherence-XY_West.pdf
Coherence-XY_West.pdf
Attachment 7: Coherence-YZ_West.pdf
Coherence-YZ_West.pdf
  171   Wed Dec 5 20:32:51 2007 tobinUpdatePSLISS dark noise
The ISS dark noise is not coming from the PD heads; the spectrum is essentially unchanged when the PD is unplugged from the ISS. Did the input opamps both get semi-fried in the same way? (They worked so well when they were first installed.) What else changed? I'm baffled. Frown
  170   Wed Dec 5 19:25:07 2007 ranaDAQCDSDMF
I made a database file on C1AUX called dmf.db. It has 9 DMF EPICS channels which are also trended
so that one can now write data to those channels from a DMF Monitor and the data will be records.

New channels:
[C1:DMF-SEIS_1]
[C1:DMF-SEIS_2]
[C1:DMF-SEIS_3]
[C1:DMF-LINE_1]
[C1:DMF-LINE_2]
[C1:DMF-LINE_3]
[C1:DMF-MC_1]
[C1:DMF-MC_2]
[C1:DMF-MC_3]

I added these to C1AUX because it doesn't do much and can be booted without having much effect.
(it controls Mech Shutters, Video, and Illuminators. It used to also do the EO Shutter but I
removed that from its startup.cmd and it will no longer load those records).
  169   Wed Dec 5 18:22:03 2007 tobinUpdatePSLISS dark noise
Attached is a plot of the dark noise spectrum of the ISS photodiodes (1) before fooling with them, (2) after replacing the 4151's with 4131's (improvement!), and (3) after replacing the cables and changing the wiring (disaster!).
Attachment 1: sense_noise.pdf
sense_noise.pdf
  168   Wed Dec 5 18:08:36 2007 AndreyUpdateASCOptical Lever laser for ETMX is installed

A new laser with \lambda=633nm has been intalled and the mirror adjusted so that the signal hits the center of the photodetector.

Output power level of that laser is 3.45 +- 0.05 mW.

Only about 0.29mW hits the photodetector.

Cable clips have been used to firmly fix the power supply cable for the laser.

See attached photopicture of the ETMX - "oplev" - optical - table.
Attachment 1: DSC_0199.JPG
DSC_0199.JPG
  167   Wed Dec 5 17:49:57 2007 tobinUpdatePSLISS
Attached is a plot of the ISS RIN with a variety of gain settings.

Unfortunately the dark noise is huge now--a result of the new cables & wiring?
Attachment 1: rin.pdf
rin.pdf
  166   Wed Dec 5 16:57:36 2007 tobinHowToComputer Scripts / ProgramsSR785 data converter on linux
I was pleased to find that the SR785 Data Viewer (including the command line conversion utility) installs and works in linux using WINE (on my laptop at least). There are some quirks, of course, but I was able to extract my data.
  165   Wed Dec 5 13:49:08 2007 albertoUpdateElectronicsRF AM PD lines monitor
In the last weeks Iíve been working on the design of an electronic board to measure directly the power of the main spectral lines on of the RF-AM photodiode from as many independent outputs. The idea is to have eventually a monitor channel in the CDS network for the power of each line.
Looking at at the spectrum from the RF-AM PD (see attached plot), there are 5 main lines:
Frequency
3 fsr = 33 195 439 Hz
4 fsr = 66 390 878 Hz
12 fsr = 132 781 756 Hz
15 fsr = 165 977 195 Hz
18 fsr = 199 172 634 Hz

Two main approaches have been proposed for the circuit depending on the way followed to isolate the lines:
1) Filters: the frequencies are separated by narrow notch filters, then a diode bridge rectifies and a low pass filter extracts the DC component.
2) Mixers: for each frequency there is a mixer driven by a copy of the correspondent modulation frequency provided by the function generators (the Marconi). The mixers automatically give the DC component of the rectified signals.
Because of the phase lags that we should compensate if we used mixers, we would prefer the first approach, if it works.
Starting with a tolerance of about 10% between the channels, the spectrum (see attachment) sets the constraint to the filterís suppression:
Filter central frequency [MHz]******Suppression within 30 Mhz [dB]
33*********************************-7-20 = -27
66**********************************7-20 = -13
133*********************************12-20 = -8
166********************************-12-20 = -32
199*********************************10-20 = 10

So far Iíve tried two kinds of designs for the filters, Butterworth (see attachment) and LC and I'm measuring transfer functions tuning the components to match the central frequency and the bandwdth of the filters with the requirements.

The frequencies weíre dealing with are rather high and several adjustments had to be done to the measurement system in order to shield the circuit from the impedance of the input and the output line (i.e., amplifier turned out to be necessary). Also, an the mixer had to be replaced to an RF one.
It seems I'm now measuring new transfer functions (which look quite different from what I've got with no amplifiers).
To be posted soon.
Attachment 1: alberto.spectrum2.png
alberto.spectrum2.png
Attachment 2: Butterworth.PNG
Butterworth.PNG
  164   Wed Dec 5 10:57:08 2007 albertoHowToComputersConnecting the GPIBto USB interface to the Dell laptop
The interface works only on one of the USB ports of the laptop (the one on the right, looking at the computer from the back).
  163   Tue Dec 4 23:16:35 2007 tobinUpdatePSLISS
I was confused to find that I could increase the ISS gain slider all the way from 15dB to 30dB without seeing much of any increase in gain in the measured open-loop transfer function. While making these swept-sine measurements, the saturation indicator almost never tripped, indicating it was seemingly happy. But then I noticed an odd thing: if I disable the test ("analog excitation") input, the saturation indicator trips immediately. I hooked up a scope to the current shunt test point (TP12). With the test input enabled, the loop closed, and the analog excitation port connected to the SR785, I see a a 5 Vpkpk, 2.55 MHz triangle wave there. It is there even if I set the SR785 excitation amplitude to zero, but it disappears if I disconnect the cable from the SR785.

I found oscillations at TP20, TP30, TP36, TP41, and TP42. Many of these are in the (unused) "outer loop" circuitry and currently lack compensation capacitors.
ELOG V3.1.3-